WHAT IS THE CORRECT STATEMENT OF FIRST LAW OF THERMODYNAMICS? WHAT IS ITS CORRECT EQUATION.


IN OUR CHEMISTRY BOOK & PHYSICS BOOK THEY HAVE GIVEN DIFFERENT EQUATIONS.


ALSO TELL WHEN WORK DONE IS POSITIVE & WHEN IS IT NEGATIVE ?


SAUMYA

Grade:11

2 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Saumya

In Physics

U=Q-W

when work is done by the system W is positive and when work is done on the system W is negative

When heat is given to system Q is positive and when heat is given by the system Q is negative

In Chemistry

U=Q+W

when work is done by the system W is negative and when work is done on the system W is positive

When heat is given to system Q is positive and when heat is given by the system Q is negative

Hence both are same with a difference in sign convention only.

 

All the best.                                                           

AKASH GOYAL

AskiitiansExpert-IIT Delhi

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

 


AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Saumya

In Physics

U=Q-W

when work is done by the system W is positive and when work is done on the system W is negative

When heat is given to system Q is positive and when heat is given by the system Q is negative

In Chemistry

U=Q+W

when work is done by the system W is negative and when work is done on the system W is positive

When heat is given to system Q is positive and when heat is given by the system Q is negative

Hence both are same with a difference in sign convention only.

 

All the best.                                                           

AKASH GOYAL

AskiitiansExpert-IIT Delhi

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..